OpenStudy (perl):

I have a question about quarternions. These are an extension of complex numbers. Given that i^2 = j^2 = k^2 = ijk=-1 show that ij = k = -ji jk = i = -kj ki = j = -ki I am having trouble with the negatives

OpenStudy (perl):

It is easy to see that given ijk = -1 ijk *k = -1 *k ij (kk) = -1 *k ij (-1) = -1 *k ij = k

OpenStudy (perl):

now prove -ji = k , thats what im having trouble with

OpenStudy (anonymous):

Are you not allowed to take right hand multiplication?

OpenStudy (perl):

this is non-commutative. but you can do right multiplication or left multiplication. and it is associative

OpenStudy (perl):

im not seeing algebraically how to prove ji = -k

OpenStudy (perl):

oh i think i see it

OpenStudy (perl):

ij = k ij i = ki i j i = j i i j i = i j -1 ji = ij

OpenStudy (anonymous):

Oh I see, well that figures, we didn't cover that topic much yet in my university so you have the following identity \(ijk=-1\) multiply by \(ji\) from the let to get \[\Large (ji)ijk = -ji \]

OpenStudy (perl):

oh thats simpler :)

OpenStudy (perl):

the terms cancel

OpenStudy (perl):

(ji) (ijk) = (ji)(-1) j(ii)jk = -ji j(-1)jk = -ji (-1)jjk = -ji (-1)(-1)k = -ji k = -ji

OpenStudy (anonymous):

it's really just a game between left hand and right hand multiplication, but not very obvious at first glance, however if I didn't contribute to any algebraic fallacy than that would turn out to be your expression \[\Large j(ii)jk=-ji \implies -jjk=-ji \implies k=-ji \]

OpenStudy (perl):

right, i also did it another way

OpenStudy (perl):

since we easily prove that ij = k , jk = i, ki = j, we can use these and argue: ij = k ij*i = k*i iji = j i*iji = i*j (ii) ji = ij - ji = ij

OpenStudy (perl):

but your way is easier, i have to do left multiplication and right multiplication by i

OpenStudy (anonymous):

Exactly, often Quaternions get introduced in Linear Algebra courses and as Matrix Multiplication to empathize on the topic of left hand and right hand multiplication. However, your way works just as fine, verifying these identities can be very tedious. So better to have them in an exercise rather than in a test (-;

OpenStudy (perl):

yes :)

OpenStudy (perl):

i was looking for something simple and you found it

OpenStudy (anonymous):

\[\checkmark \]

OpenStudy (perl):

simple and elegant

OpenStudy (perl):

i gave up too quickly :/

OpenStudy (perl):

thanks again

OpenStudy (anonymous):

You're very welcome :-)

OpenStudy (perl):

also we have to be careful when we do imaginary or quarternions, they are not the same as those coordinate basis vectors, like in calculus. r= ai + bj + ck. this is not the same i as z = a + bi , or h = a + bi + cj + ek (quarternions)

OpenStudy (perl):

complex numbers is a + bi. , and i think theres also octonians , but im not sure

OpenStudy (perl):

octonians are not even associative

OpenStudy (perl):

so they are completely different, but quarternions find some interesting applications

OpenStudy (perl):

this is interesting you can define quarternion multiplication using ordered pairs http://en.wikipedia.org/wiki/Cayley%E2%80%93wingspanson_construction

OpenStudy (anonymous):

Exactly, it is important to realize that \(i,j,k\) are not complex numbers in the manner of Quaternions.